OpenStudy (anonymous):

f(x) = 3x + 2; g(x) = 3x - 5 Find f/g.

zepdrix (zepdrix):

\[\large\rm \color{royalblue}{f(x)=3x+2},\qquad \color{orangered}{g(x)=3x-5}\] Then,\[\large\rm \frac{\color{royalblue}{f(x)}}{\color{orangered}{g(x)}}\]just plug in the functions, ya? :)

OpenStudy (anonymous):

3x+2/3x-5?

OpenStudy (anonymous):

then what would the domain be?

zepdrix (zepdrix):

Yes, (3x+2)/(3x-5) The domain is generally going to be all real numbers. But since we have a fraction, we might have a restriction at a location.

zepdrix (zepdrix):

Recall that in the land of math, we can not divide by 0. \(\large\rm \frac{1}{0}\) <-- This is not a number, this is undefined. This is bad. So what we'll do is, we'll `let` our denominator equal 0 like this. And see if we can find these "bad" values.

zepdrix (zepdrix):

\[\large\rm 3x-5=0\]This should give us the "bad" value for x.

OpenStudy (anonymous):

5/3?

zepdrix (zepdrix):

Good good good. This is the bad value. Which means our domain consists of every x `except` x=5/3

OpenStudy (anonymous):

so would the full answer be (f/g)(x) = Quantity three x plus two divided by three x minus five.; domain {x|x ≠ Five over three. }

OpenStudy (anonymous):

Can I ask one more?

zepdrix (zepdrix):

so would the full answer be (f/g)(x) = Quantity three x plus two divided by `the quantity` three x minus five.; domain {x|x ≠ Five over three. } I would probably add this in, so it's clear that the bottom is an entire quantity, not just a 3x denominator.

zepdrix (zepdrix):

Sure

OpenStudy (anonymous):

f(x) = 7x + 7, g(x) = 6x2 Find (fg)(x).

OpenStudy (anonymous):

ok thanks:)

zepdrix (zepdrix):

\[\large\rm \color{royalblue}{f(x) = 7x + 7},\qquad\qquad \color{orangered}{g(x) = 6x^2}\] \[\large\rm (fg)(x)=\color{royalblue}{f(x)}\color{orangered}{g(x)}\]\[\large\rm =\color{royalblue}{(7x+7)}\color{orangered}{6x^2}\]So here I guess you'll have to do some distributing.

OpenStudy (anonymous):

so 42x^3+42x?

OpenStudy (anonymous):

?

zepdrix (zepdrix):

42x^3+42x^2 ya?